Avillage has total of 100 under five children. the coverage with measles vaccine in this age group is 60%. following the occurrence of ameasles case in achild after avisit outside 26 children developed measles. the secondary attack rate of measles is:-
Answer #1

Ans) 66%

Total No. of children = 100

No. of immunized = 60

No. of un-immunized = 40

Primary case = 1 ( case in achild after avisit outside 26 children developed measles. )

secondary attack rate = 26/(40-1) = 26/39 x 100 = 66%